A metal cylinder has a mass of 100.0 grams. If it measures 4.0 in high and 1.50 in wide, what is its density (ρ) in g/cm3 if 1.0 in=2.54 cm?
1
Expert's answer
2020-10-05T10:53:08-0400
"\rho=\frac{m}{\pi r^2h}\\\rho=\frac{100}{\pi (0.5\cdot1.5\cdot 2.54)^2(4\cdot 2.54)}\\\rho=0.863\frac{g}{cm^3}"